a certain confidence interval is
8.95 to 10.25
Find the sample mean and margin of error.

Answer by jim_thompson5910(35256)

To find the sample mean (xbar) of any confidence interval (L,U), we use the formula below


xbar = (U + L)/2


Plug in the given values L = 8.95 and U = 10.25 and then evaluate


xbar = (U + L)/2


xbar = (10.25 + 8.95)/2


xbar = 9.6


So the sample mean is 9.6


-------------------------


To find the margin of error (ME) of any confidence interval (L,U), we use the formula below


ME = (U - L)/2


Plug in the given values L = 8.95 and U = 10.25 and then evaluate


ME = (U - L)/2


ME = (10.25 - 8.95)/2


ME = 0.65


The margin of error is 0.65